Given that 3 + i is a root of the equation z^3 − 3z^2 − 8z + 30= 0, find the remaining roots.
1
Expert's answer
2020-05-19T08:58:49-0400

"z^3 \u2212 3z^2 \u2212 8z + 30= 0"

"(z+3)(z^2 \u2212 6z + 10)= 0"

"z+3=0" "z^2-6z+10=0"

"z^2-6z+10=0"

"D= (-6)^2 - 4*1*10=-4"

"z_1=\frac{6+\sqrt{D}}{2*1}=\frac{6+2i}{2}=3+i"

"z_2=\frac{6-\sqrt{D}}{2*1}=\frac{6-2i}{2}=3-i"


"z+3=0"

"z_3=-3"


Answer:

"z_1=3+i, z_2=3-i, z_3=-3"